# Mind the Gap: Removing the Discretization Gap in Differentiable Logic Gate Networks

### v1

@@ 5 Empirical Evaluations
Our empirical evaluations focus on CIFAR-10, as the MNIST-like datasets (MNIST, FashionMNIST, etc.) have low discretization gaps (cf. Appendix I). CIFAR-10 and MNIST were also the datasets Petersen et al. [1, 2] mostly focused on. Due to constrained resources, we limit experiments by default to 48 GPU hours. We use the hyperparameters from [1, 2] whenever possible rather than tuning the parameters, such as learning rate, ourselves. Section H.1 contains all the default parameters.
@@
Figure 3 shows the test accuracy as a function of training iteration for an LGN of depth 12 and width $256$k on CIFAR-10; these are the default parameters unless stated otherwise. To quantify the discretization gap, we take the absolute difference between the discretized and soft network accuracy as shown on the right in Figure 3. Gumbel LGN converges much faster than the Differentiable LGN, with virtually no discretization gap. We also measure the runtime and see that Gumbel increases the runtime by roughly $5\%$ per iteration (cf. Appendix K). Combined with runtime results from Table 4 in Appendix K, Gumbel LGN converges [^2] For this, we match Gumbel LGNs’ discrete accuracy with Differentiable LGNs’ maximum discrete accuracy. 4.75×4.75$\times$4.75 × faster in iterations, making Gumbel LGNs 4.5×4.5$\times$4.5 × faster in wall-clock time to train. Note that the Differentiable LGN still improves after 48 hours.
@@
On the right of Figure 3, we see the discretization gap for models of various depths for Differentiable LGNs and Gumbel LGNs. As the model depth increases, the expressive power of the networks theoretically increases. The Differentiable LGNs experience bigger discretization gaps as the depth increases, while our Gumbel LGNs are stable across depths. Hence, Gumbel LGNs do not struggle to converge as the networks are made deeper. Both methods experience a shift in when the maximum gap occurs. This is expected since increasing the model size usually delays when the accuracy plateaus, i.e., more iterations are needed to converge.
Figure 4: Straight-through (ST) estimator ablation. The Gumbel LGNs uses hard gate choices in the forward pass as shown in (2) called the ST estimator. On the left, we show the test accuracy over training iterations; on the right, we show the discretization gap. Gumbel LGNs with ST estimator converge slightly slower in test accuracy, but the discretization gap is smaller.
Figure 5: Ablation over the temperature $\tau$ for Gumbel LGNs. Left: Estimated Hessian trace using Hutchinson’s method. The trace shrinks as $\tau$ decreases, indicating fewer large positive eigenvalues, thus suggesting a flatter loss surface, which may reduce the risk of loss increases when discretizing parameters. Right: Test accuracy for the $\tau$-values. We see a goldilocks zone for the temperature, as if $\tau$ is large ($>1$) or small ($<0.1$), then the network converges much slower. In the zoomed-in view, we plot the non-discretized view as dashed lines and see that these are similar to the discretized values, i.e., the discretization gap is low for all. Except for $\tau=0.01$ and $\tau=2$, there is only a small variation in the final accuracy, as seen in Table 1.
@@
To better understand the source of gap reduction achieved by the Gumbel-Softmax trick, we perform an ablation to isolate the contribution of the ST estimator. As discussed previously, Gumbel-Softmax combines (i) ST estimation as seen in (2) and (ii) implicit smoothing via Gumbel noise. By disabling the ST path, we aim to identify whether gap reduction primarily stems from the discrete gradient approximation or the added stochasticity. The setup without ST corresponds to Differentiable LGNs with noisy logits and is denoted as Soft Gumbel.
In Figure 4, we see that imputing Gumbel noise alone impacts both convergence and discretization compared to Differentiable LGNs. However, we observe that including the ST estimator delays convergence for a fixed $\tau$, but further reduces the discretization gap.
@@
As we showed in Section 4, loss minimization for Gumbel LGNs implicitly reduces the curvature of minima by minimizing the trace of the Hessian. The impact of the trace compared to the loss depends on the temperature parameter $\tau$. Thus, we perform an ablation study on the effect of $\tau$, estimate the trace of the Hessian, and visualize the loss landscape.
@@
We evaluate how varying the temperature $\tau$ affects optimization dynamics. Recall that higher $\tau$ reduces the degree of implicit smoothing, potentially leading to sharper minima and slower convergence. In our experiment, we test $\tau\in[0.01,2.0]$. We show the results in Tables 1 and 5 where we observe a goldilocks zone for the temperature; if $\tau$ is large ($>1$) or small ($<0.1$), then the network converges much slower. However, higher temperatures such as $1$ seem to converge to slightly better solutions. Although the difference is minor, $<0.5\%$ when $\tau=0.25$ goes to $\tau=1$.
Table 1: Maximum and final test accuracy for the tested $\tau$ values. The iterations column indicates the number of training iterations ($\cdot 10^{3}$) required to be within $1\%$ of the maximum accuracy. We see that with a medium value of $\tau\approx 0.25$ the network converges much faster than for high values $>1$.
@@
Figure 6: Visualization of loss landscapes. Left: Loss landscape of a Differentiable LGN. We see that the landscape is overall noisy. Right: Loss landscape of a Gumbel LGN with $\tau=1.0$. We observe a much smoother loss landscape compared to the Differentiable LGNs.
@@
The Hessian scales quadratically with model size, so direct computations are infeasible for our networks with millions of parameters. Still, we can use iterative methods to approximate the trace, etc. [51, 52, 53, 54, 55]. We approximate the trace using Hutchinson’s method with 200 Rademacher random vectors, where each coordinate is independently sampled from $\{-1,+1\}$ with equal probability [54]. Full experimental details are provided in Appendix F.
As shown in Figure 5, decreasing the $\tau$-parameter reduces the estimated Hessian trace. This aligns with the theoretical insights from Section 4; lower $\tau$ values place greater weight on trace reduction. The trace is negative, this is expected: stochastic gradient noise in overparameterized networks tends to systematically lower the expected Hessian trace, biasing solutions toward flatter regions of the loss landscape that may have negative trace values [56, 57, 58]. Large negative eigenvalues often vanish, the trace remains influenced by many small eigenvalues, resulting in a negative overall trace.
@@
To qualitatively assess the loss landscape curvature, we project the high-dimensional parameter space onto two-dimensional subspaces. Following Li et al. [59], we select random directions and interpolate the loss surface along these axes, providing insight into the optimization landscape’s geometry around learned solutions. Visualization details are in Appendix G. Figure 6 shows that Gumbel LGN has a visually smoother loss surface.
@@
Petersen et al. [1] noted that neurons collapse to single gates, but they did not investigate the extent of the collapse. We examine this through neuron entropy by sampling 100k newly initialized neurons and computing the 95% interval. We also estimate expected entropy theoretically (see Appendix J), giving us a baseline distribution for neurons that have not learned. As neurons collapse, their entropy converges to 0. Figure 7 shows that many early Differentiable LGN layers do not collapse, while Gumbel LGN neurons converge with entropies near 0. Defining unused gates as those with entropy above the 2.5%-percentile threshold, Gumbel LGNs have 0.00% and Differentiable LGNs have 49.81% unused gates, representing a 100.00% reduction by Gumbel.
Figure 7: Entropy distribution for neurons in each layer for trained Differentiable LGNs and Gumbel LGNs models. The dashed black line indicates the expected entropy (cf. Appendix J) of neurons before training, and the shaded region is the 95% interval computed by sampling. We see that almost all the neurons in Gumbel LGNs have converged, while neurons in early layers of Differentiable LGNs still have high entropy.
@@ 6 Limitations
While Gumbel LGNs demonstrate significant improvements for deeper networks, limitations do remain. Our evaluation focuses primarily on CIFAR-10, with limited exploration of more complex datasets or problems with many output classes. We plan to thoroughly evaluate the methods on datasets like CIFAR-100 and ImageNet32 [60, 61].
The temperature parameter $\tau$ requires tuning to balance convergence speed and accuracy. Finally, our theoretical analysis connects Gumbel noise to Hessian trace minimization under simplifications, but a comprehensive theoretical treatment of the discretization gap remains an open challenge.
@@ 7 Conclusion
We introduced Gumbel logic gate networks (Gumbel LGNs), addressing two critical limitations of Differentiable LGNs: slow convergence during training and a large discretization gap between training and inference. Our theoretical analysis shows that Gumbel noise during gate selection promotes flatter minima by implicitly minimizing the Hessian trace, reducing sensitivity to parameter discretization. Experiments on CIFAR-10 demonstrate that Gumbel LGNs converge 4.5×4.5$\times$4.5 × faster in wall-clock time than Differentiable LGNs while reducing the discretization gap by $98\%$ and achieving $100.0\%$ improvement in neuron utilization. These advantages become more pronounced with depth, indicating favorable scaling properties. Our improvements are dataset and architecture-independent, and several promising directions remain for future exploration: extending Differentiable LGNs to convolutional architectures, validating on more complex datasets like CIFAR-100 and ImageNet32, exploring hardware-specific optimizations, and investigating adaptive temperature scheduling.

### v2

@@ 5 Empirical Evaluations
Our empirical evaluations focus on CIFAR-10 and CIFAR-100. Due to constrained resources, we limit experiments by default to 48 GPU hours. To ensure a fair comparison, we use the hyperparameters from [7] whenever possible rather than tuning the parameters, such as learning rate, ourselves. Section˜H.1 contains all the default parameters.
@@
Figure˜3 shows the test accuracy as a function of training iteration for an LGN of depth 12 and width 256256k on CIFAR-10; these are the default parameters unless stated otherwise. To quantify the discretization gap, we take the absolute difference between the discretized and soft network accuracy as shown on the right in Figure˜3. Gumbel LGN converges much faster than the DLGNs, with virtually no discretization gap. We also measure the runtime and see that Gumbel increases the runtime by roughly $5\%$ per iteration (cf. Appendix˜K). Combined with runtime results from Table˜4 in Appendix˜K, Gumbel LGN converges [^1] For this, we match Gumbel LGNs’ discrete accuracy with DLGNs’ maximum discrete accuracy. $4.75\times$ faster in iterations, making Gumbel LGNs $4.5\times$ faster in wall-clock time to train. Note that the Differentiable LGN still improves after 48 hours.
Figure 3: Performance of Gumbel LGNs and DLGNs on CIFAR-10. (Left) Test accuracy (width of 256256k, depth of 12). (Right) Discretization gap for various depths. DLGNs experience larger gaps and slower reduction as the depth increases. In contrast, Gumbel LGNs have consistently low gaps and fast reduction as the network depth increases.
We perform an additional evaluation on CIFAR-100 using the same settings. On Figure 4, we see that GLGNs already converge at around 400K iterations, whereas DLGNs do not converge. This is consistent with the results, we observe for CIFAR-10.
@@
On the right of Figure˜3, we see the discretization gap for models of various depths for DLGNs and Gumbel LGNs. As the model depth increases, the expressive power of the networks theoretically increases. The DLGNs experience bigger discretization gaps as the depth increases, while our Gumbel LGNs are stable across depths. Hence, Gumbel LGNs do not struggle to converge as the networks are made deeper. Both methods experience a shift in when the maximum gap occurs. This is expected since increasing the model size usually delays when the accuracy plateaus, i.e., more iterations are needed to converge.
Figure 4: Performance of Gumbel LGNs and DLGNs on CIFAR-100. (Left) Test accuracy (width of 256256k, depth of 12). (Right) Discretization gap for various depths. DLGNs experience larger gaps and slower reduction as the depth increases. In contrast, Gumbel LGNs have consistently low gaps and fast reduction as the network depth increases.
Shallow Network
We fix the depth of the network at 6 and increase the width to 2048K. On Figure 5, we see that. Interestingly, the accuracy of the soft DLGN converges within 50K iterations. However, the discretized version does not seem to improve with continued training. We note that the final accuracy of GLGNs is higher than DLGNs. This suggests increasing the number of neurons through the width also contributes to the discretization gap.
Figure 5: Test accuracy on CIFAR-10 for a shallow, wide network (width 2048k, depth 6).
Figure 6: Straight-through (ST) estimator ablation (width of 256k, depth 12). The Gumbel LGNs uses hard gate choices in the forward pass as shown in (2) called the ST estimator. On the left, we show the test accuracy over training iterations; on the right, we show the discretization gap. Gumbel LGNs with ST estimator converge slightly slower in test accuracy, but the discretization gap is smaller.
Figure 7: Ablation over the temperature $\tau$ for Gumbel LGNs. Left) Estimated Hessian trace using Hutchinson’s method. The trace shrinks as $\tau$ decreases, indicating fewer large positive eigenvalues, thus suggesting a flatter loss surface, which may reduce the risk of loss increases when discretizing parameters. Right) Test accuracy for the $\tau$-values. We see a goldilocks zone for the temperature, as if $\tau$ is large (>1>1) or small (<0.1<0.1), then the network converges much slower. In the zoomed-in view, we plot the non-discretized view as dashed lines and see that these are similar to the discretized values, i.e., the discretization gap is low for all. Except for $\tau=0.01$ and $\tau=2$, there is only a small variation in the final accuracy, as seen in Table˜1.
@@
To better understand the source of gap reduction achieved by the Gumbel-Softmax trick, we perform an ablation to isolate the contribution of the ST estimator. As discussed previously, Gumbel-Softmax combines (i) ST estimation as seen in (2) and (ii) implicit smoothing via Gumbel noise. By disabling the ST path, we aim to identify whether gap reduction primarily stems from the discrete gradient approximation or the added stochasticity. The setup without ST corresponds to DLGNs with noisy logits and is denoted as Soft Gumbel.
In Figure˜6, we see that imputing Gumbel noise alone impacts both convergence and discretization compared to DLGNs. However, we observe that including the ST estimator delays convergence for a fixed $\tau$, but further reduces the discretization gap.
@@
As we showed in Section˜4, loss minimization for Gumbel LGNs implicitly reduces the curvature of minima by minimizing the trace of the Hessian. The impact of the trace compared to the loss depends on the temperature parameter $\tau$. Thus, we perform an ablation study on the effect of $\tau$, estimate the trace of the Hessian, and visualize the loss landscape.
@@
We evaluate how varying the temperature $\tau$ affects optimization dynamics. Recall that higher $\tau$ reduces the degree of implicit smoothing, potentially leading to sharper minima and slower convergence. In our experiment, we test $\tau\in[0.01,2.0]$. We show the results in Tables˜1 and 7 where we observe a goldilocks zone for the temperature; if $\tau$ is large (>1>1) or small (<0.1<0.1), then the network converges much slower. However, higher temperatures such as $1$ seem to converge to slightly better solutions. Although the difference is minor, $0.5\%$ when $\tau=0.25$ goes to $\tau=1$.
Table 1: Maximum and final test accuracy for the tested $\tau$ values. The iterations column indicates the number of training iterations ($\cdot 10^{3}$) required to be within $1\%$ of the maximum accuracy. We see that with a medium value of $\tau\approx 0.25$ the network converges much faster than for high values >1>1.
@@
Figure 8: Visualization of loss landscapes. Left: Loss landscape of a Differentiable LGN. We see that the landscape is overall noisy. Right: Loss landscape of a Gumbel LGN with $\tau=1.0$. We observe a much smoother loss landscape compared to the DLGNs.
@@
The Hessian scales quadratically with model size, so direct computations are infeasible for our networks with millions of parameters. Still, we can use iterative methods to approximate the trace, etc. [64, 65, 66, 67, 68]. We approximate the trace using Hutchinson’s method with 200 Rademacher random vectors, where each coordinate is independently sampled from $\{-1,+1\}$ with equal probability [67]. Full experimental details are provided in Appendix˜F.
As shown in Figure˜7, decreasing the $\tau$-parameter reduces the estimated Hessian trace. This aligns with the theoretical insights from Section˜4; lower $\tau$ values place greater weight on trace reduction. The trace is negative, this is expected: stochastic gradient noise in overparameterized networks tends to systematically lower the expected Hessian trace, biasing solutions toward flatter regions of the loss landscape that may have negative trace values [69, 70, 71]. Large negative eigenvalues often vanish, the trace remains influenced by many small eigenvalues, resulting in a negative overall trace.
@@
To qualitatively assess the loss landscape curvature, we project the high-dimensional parameter space onto two-dimensional subspaces. Following Li et al. [72], we select random directions and interpolate the loss surface along these axes, providing insight into the optimization landscape’s geometry around learned solutions. Visualization details are in Appendix˜G. Figure˜8 shows that Gumbel LGN has a visually smoother loss surface.
@@
Petersen et al. [7] noted that neurons collapse to single gates, but they did not investigate the extent of the collapse. We examine this through neuron entropy by sampling 100k newly initialized neurons and computing the 95% interval. We also estimate expected entropy theoretically (see Appendix˜J), giving us a baseline distribution for neurons that have not learned. As neurons collapse, their entropy converges to 0. Figure˜9 shows that many early Differentiable LGN layers do not collapse, while Gumbel LGN neurons converge with entropies near 0. Defining unused gates as those with entropy above the 2.5%-percentile threshold, Gumbel LGNs have 0.00% and DLGNs have 49.81% unused gates, representing a 100.00% reduction by Gumbel.
Figure 9: Entropy distribution for neurons in each layer for trained DLGNs and Gumbel LGNs models. The dashed black line indicates the expected entropy (cf. Appendix˜J) of neurons before training, and the shaded region is the 95% interval computed by sampling. We see that almost all the neurons in Gumbel LGNs have converged, while neurons in early layers of DLGNs still have high entropy.
@@ 6 Limitations
While Gumbel LGNs demonstrate significant improvements for deeper networks, limitations do remain. Our evaluation focuses primarily on CIFAR-10 and CIFAR-100, with limited exploration of more complex datasets. The temperature parameter $\tau$ requires tuning to balance convergence speed and accuracy. Finally, our theoretical analysis connects Gumbel noise to Hessian trace minimization under simplifications, but a comprehensive theoretical treatment of the discretization gap remains an open challenge. Moreover, a comprehensive analysis of width scaling and the interplay between width and depth is left for future work.
@@ 7 Conclusion
We introduced Gumbel Logic Gate Networks (Gumbel LGNs), addressing two critical limitations of DLGNs: slow convergence during training and a large discretization gap. Our theoretical analysis shows that Gumbel noise during gate selection promotes flatter minima by implicitly minimizing the Hessian trace, reducing sensitivity to parameter discretization. Experiments on CIFAR-100 and CIFAR-10 demonstrate that Gumbel LGNs converge up to $4.5\times$ faster in wall-clock time than DLGNs while reducing the discretization gap by $98\%$ and achieving $100.0\%$ improvement in neuron utilization. These advantages become more pronounced with depth, indicating favorable scaling properties. Our improvements are dataset and architecture-independent, and several promising directions remain for future exploration, e.g. adaptive temperature scheduling.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification:
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification:
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: Proof in Appendix˜B.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: See, for instance, Sections˜H.1 and G.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: CIFAR-10 is an open-source dataset and we provide the codebase.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: See Section˜H.1.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [No]
  34. Justification: We were not able to rerun the experiments under multiple seeds due to compute constraints.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: See Appendix˜L.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification:
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [N/A]
  49. Justification: We conduct foundational research.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: To the best of our knowledge, there are no risks posed by the release of our code.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: We only used open-sourced models and data.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: We provide instructions for the codebase in the zip file.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: NA
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: NA
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: NA
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.